OpenStudy (anonymous):

Medal will be awarded) The radius of a cylinder is 5.4 X 10^5 cm. The height of the cylinder is 2.5 X 10^3 cm. What is the volume of the cylinder? (Hint: V=Pi r^2 h) (I'm on multiplication properties of exponents. How do I start?)

OpenStudy (anonymous):

I don't know how to start!

OpenStudy (kropot72):

\[V=\pi \times r ^{2}\times h\] where r is the radius and h is the height. Now you need to substitute the given values of r and h and calculate the answer in cubic centimeters.

OpenStudy (anonymous):

Alright. One sec.

OpenStudy (anonymous):

what do I do? I have \[\pi(29.16 X 10^{12} 2.5 X 10^{3}\]

OpenStudy (kropot72):

Nearly correct. Just one thing needs changing. When dealing with exponents consider this rule: \[(a ^{b})^{c}=a ^{bc}\] So what is \[(10^{5})^{2}=?\]

OpenStudy (anonymous):

10^10?

OpenStudy (anonymous):

I'm asking if I need to multiply the 29.16 and 2.5 and then mult. that by pi.

OpenStudy (anonymous):

Like do I get 228.906 X 10^36 and then I need to make it 2.28906 X 10^34?

OpenStudy (kropot72):

Good work! So your calculation now becomes: \[V=\pi \times 29.16\times 10^{10}\times 2.5\times 10^{3}=?\]

OpenStudy (anonymous):

So is the answer \[V= 2.28906 X 10^{34}\]

OpenStudy (anonymous):

Do you mind if I ask you a bit more questions after this?

OpenStudy (kropot72):

You still need to work on this question. If I rearrange the equation it should make the process clearer: \[V=\pi \times 29.16\times 2.5\times 10^{10}\times 10^{3}=?\] Hint: When multiplying the two terms that are powers of 10 remember the rule: \[a ^{b}\times a ^{c}=a ^{b+c}\]

OpenStudy (anonymous):

228.906 X 10^30 (or 10^13?)

OpenStudy (kropot72):

I get the following: \[V=229.022\times 10^{13}\ cm ^{3}\]

OpenStudy (anonymous):

How did you get 229.022? I got 228.906 (i'm using 3.14 for pi)

OpenStudy (kropot72):

The difference comes about because I am using the value for pi that is programed into my calculator. If you are told to use 3.14 for pi then you are correct.

OpenStudy (anonymous):

Alright. Can I ask you some more questions? Or do you need to leave?

OpenStudy (anonymous):

Thanks for all your help!

OpenStudy (kropot72):

Remember these two rules for exponents: \[(a ^{b})^{c}=a ^{bc}\] \[a ^{b}\times a ^{c}=a ^{b+c}\] Sorry but I must log out now. Please post more questions separately and others will help :)